# Critical states exhibit invariance in both position and momentum spaces

### v1

@@ Abstract
The critical states of disordered systems are intriguing subjects within the realm of condensed matter physics and complex systems. These states manifest in materials where disorder plays a significant role, and are distinguished by their multifractal structure and self-similarity. However, accurately characterizing critical states continues to pose a significant challenge. In this study, we argue that critical states exhibit a certain invariance in both position and momentum spaces, leading to their delocalization in both domains. More specifically, it is expected that typical physical quantities characterizing critical states, such as the inverse participation ratio and information entropy, should exhibit invariance in both position space and momentum space. Subsequent numerical simulations validate the correctness of this invariance, thereby establishing a robust foundation for future experimental validation of critical states.
@@ I Introduction
Recently, significant advancements have been achieved in the investigation of critical states of wave functions, specifically focusing on critical states of quantum disordered systems Liu-self-duality ; Li1 ; Li2 ; You ; g1 ; g2 ; zhou ; Biddle ; Zhang ; Chen ; q1 ; q2 ; q3 ; q4 ; q5 ; q6 ; q7 ; q8 ; Yao1 ; Yao2 ; cai ; Slager ; SSHQ ; gg3 ; gg4 ; gg5 ; gg6 ; gg7 ; gg8 ; gg0 ; gg9 . Novel concepts and phenomena such as anomalous mobility edges Liu-Anomalous , the utilization of renormalization groups for critical states RG1 ; RG2 , real eigenvalues determined through recursion of eigenstates Liu-Real1 ; Liu-Real2 , critical states induced by coupling of two chains twochain , and critical states of two-dimensional quasicrystals Duncan are continuously emerging. Nevertheless, the primary challenge in comprehending critical states lies in determining the conditions under which they arise and accurately characterizing them.
@@
Figure 1: (Color online) (a) and (b) demonstrate an application of Bohr’s complementarity principle, showing that a localized state in position space must be delocalized in momentum space, and vice versa for an extended state. (c) and (d) illustrate that critical states can simultaneously exhibit delocalization, multifractality, and self-similarity in both position space and momentum space.
@@
A natural question arises: can the invariance of the critical state extend beyond the Lyapunov exponent to include other physical quantities in both position and momentum spaces? This study demonstrates that the invariance of the critical state is universal, encompassing various physical quantities such as the inverse participation ratio and information entropy.
@@ II Analytical derivation
The Inverse Participation Ratio (IPR) is a physical quantity IPR ; IPR1 ; IPR2 commonly utilized in the fields of quantum mechanics and condensed matter physics to analyze the localization of wave functions Li3 ; Li4 ; gg1 ; gg2 , particularly within disordered systems. It serves as a quantitative measure of the extent to which a wave function is distributed across space versus its degree of localization. A large IPR indicates high localization, with the majority of probability density concentrated in a small region, while a smaller IPR reflects a more uniform distribution of probability density across multiple sites.
Information entropy $S$, commonly known as Shannon entropy Shannon in honor of Claude Shannon, is a metric for the unpredictability associated with a system. Initially introduced in the context of information theory, it has since been widely applied, particularly in the examination of quantum complex systems IE1 . Information entropy has been utilized to comprehend and quantify the level of unpredictability and randomness in disordered systems IE2 ; IE3 .
@@ III.2 Quasiperiodic-Nonlinear-Eigenproblem model
Figure 2: (Color online) (a), (b) and (c) illustrate the comparison of critical states, IPR and $S$ in position space and momentum space for the AAH model, respectively. (d), (e) and (f) demonstrate the comparison of critical states, IPR and $S$ in position space and momentum space for the QNE model, respectively. Notably, critical states $|\psi|$ and $|\phi|$ exhibit simultaneous delocalization, multifractality, and self-similarity in both position space and momentum space. In the AAH model, IPR and $S$ are only equal between position space and momentum space at the phase transition point $V=2$, whereas in the QNE model, these physical quantities are equal within a large parameter range of $0<V\leq 2$.
@@ IV Summary
In conclusion, we have addressed a long-standing unresolved issue regarding the definition of a critical state and its accurate characterization. Our findings emphasize the universality of critical state invariance across various domains in physics.

### v2

@@ Abstract
The critical states of disordered systems are intriguing subjects within the realm of condensed matter physics and complex systems. These states manifest in materials where disorder plays a significant role, and are distinguished by their multifractal structure and self-similarity. However, accurately characterizing critical states continues to pose a significant challenge. In this work, we propose an Ansatz for the critical state, and argue that critical states exhibit a certain invariance in both position and momentum spaces, leading to their delocalization in both domains. More specifically, it is expected that typical physical quantities characterizing critical states, such as the inverse participation ratio and information entropy, should exhibit invariance in both position space and momentum space. Subsequent numerical simulations validate the correctness of this invariance, thereby establishing a robust foundation for future experimental validation of critical states.
@@ I Introduction
Recently, significant advancements have been achieved in the investigation of critical states of quantum disordered systems, specifically focusing on multifractal critical wave functions Liu-self-duality ; Li1 ; Li2 ; You ; g1 ; g2 ; zhou ; Biddle ; Zhang ; Chen ; q1 ; q2 ; q3 ; q4 ; gg10 ; gg11 ; gg12 ; q5 ; q6 ; q7 ; q8 ; Yao1 ; Yao2 ; cai ; Slager ; SSHQ ; gg3 ; Jiang1 ; Jiang2 ; gg5 ; gg6 ; gg7 ; gg8 ; gg0 ; gg9 . Novel concepts and phenomena such as anomalous mobility edges Liu-Anomalous , the utilization of renormalization groups for critical states RG1 ; RG2 , real eigenvalues determined through recursion of eigenstates Liu-Real1 ; Liu-Real2 , critical states induced by coupling of two chains twochain , and critical states of two-dimensional quasicrystals Duncan are continuously emerging. Nevertheless, the primary challenge in comprehending critical states lies in establishing an explicit definition of the critical state and accurately characterizing them.
@@
Figure 1: (Color online) Three typical wave functions of disordered systems in dual spaces. (a) and (b) demonstrate that a localized state in position space must be delocalized in momentum space, and vice versa for an extended state. (c) and (d) illustrate that critical states can simultaneously exhibit delocalization, multifractality, and self-similarity in both position space and momentum space.
@@
Starting from a precise definition, it is widely known that the wave function of the extended state is like $\exp(i\tilde{k}n)$, while the wave function of the localized state is like $\exp(\tilde{k}n)$, where $\tilde{k}$ is a real number. Contrastively, so far we have not obtained a closed-form solution for the critical state, not even an approximate expression. Based on the delocalization, multifractality and self-similarity of critical states, we propose a hypothetical solution, that is, the critical state is a power-law modulated Weierstrass-like function,
$\psi_{\text{critical}}\sim\sum_{k=0}^{\infty}k^{c}a^{k}\cos(b^{k}\pi n),$ (1)
where $10<a<$ 1, $b$ is a positive number, and the index $c<0$ is a constant. Thus, we provide a probable expression for the critical state.
A natural question arises: can the invariance of the critical state extend beyond the Lyapunov exponent to include other physical quantities in both position and momentum spaces? This study demonstrates that the invariance of the critical state is universal, encompassing commonly utilized observables in experiments such as the inverse participation ratio and information entropy.
@@ II Analytical derivation
The Inverse Participation Ratio IPR ; IPR1 ; IPR2 ; IPR3 (IPR) and the associated participation entropy Fan1 ; Fan2 ; Fan3 ; Optical1 ; Optical2 are commonly used physical quantities in experiments to analyze the localization of wave functions Li3 ; Li4 ; gg1 ; gg2 . A large IPR indicates high localization, with the majority of probability density concentrated in a small region, while a smaller IPR reflects a more uniform distribution of probability density across multiple sites.
Another typical physical quantity, information entropy $S$, commonly known as Shannon entropy Shannon , is a metric for the unpredictability associated with a system. Initially introduced in the context of information theory, it has since been widely applied, particularly in the examination of quantum disordered systems multifractal1 ; multifractal2 ; IE1 . Utilized to quantify the level of unpredictability and uncertainty IE2 ; IE3 , a large $S$ indicates a high unpredictability, corresponding to extended states; while a low $S$ reflects a high predictability, corresponding to localized states.
@@ III.2 Quasiperiodic-Nonlinear-Eigenproblem model
Figure 2: (Color online) (a), (b) and (c) illustrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the AAH model, respectively. (d), (e) and (f) demonstrate the comparison of wave functions, IPR and $S$ in position and momentum spaces for the QNE model, respectively. Notably, critical states $|\psi|$ and $|\phi|$ exhibit simultaneous delocalization, multifractality, and self-similarity in both position space and momentum space. In the AAH model, IPR and $S$ are only equal between position space and momentum space at the phase transition point $V=2$, whereas in the QNE model, these physical quantities are equal within a large parameter range of $0<V\leq 2$.
@@ IV Summary
In conclusion, we propose an Ansatz for the critical state, and address a long-standing unresolved issue regarding the definition of a critical state and its accurate characterization. Our findings emphasize the universality of critical state invariance across various domains in physics.